Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours. In the most recent month, 195,000 items were shipped to customers using 8,600 direct labor-hours. The company incurred a total of $30,530 in variable overhead costs. According to the company’s standards, 0.04 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.60 per direct labor-hour. Required: 1. What is the standard labor-hours allowed (SH) to ship 195,000 items to customers? 2. What is the standard variable overhead cost allowed (SH × SR) to ship 195,000 items to customers? 3. What is the variable overhead spending variance? 4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

Standard hours allowed = 195,000*0.04

= 7800

Standard cost allowed = 7800*3.60

= 28080

Voh spending variance = 28080 - 30530

= 2450 U

Voh rate variance = (SR-AR) *AH

= (3.60-30530/8600)*8600

= 430 F

Voh efficiency variance = (SH-AH) * SR

= (7800 - 8600) * 3.60

= 2880 U